### v1

# Population-Proportional Preference Learning from Human Feedback: An Axiomatic Approach

Abstract
Conventional preference learning methods often prioritize opinions held more widely when aggregating preferences from multiple evaluators. This may result in policies that are biased in favor of some types of opinions or groups. The objective of this paper is to develop a novel preference learning framework capable of aligning aggregate opinions and policies proportionally with the true population distribution of evaluator preferences. Our approach infers the feasible set of evaluator population distributions directly from pairwise comparison data. Using these estimates, the algorithm constructs a policy that satisfies foundational axioms from social choice theory, namely monotonicity and Pareto efficiency, as well as our newly-introduced axioms of population-proportional representation and population-bounded robustness. We propose a soft-max relaxation method that smoothly trade-offs population-proportional representation with the selection of the Condorcet winner (which beats all other options in pairwise comparisons). Finally, we validate the effectiveness and scalability of our approach through experiments on both tabular recommendation tasks and large-scale language model alignment.
1 Introduction
Aligning artificial intelligence (AI) systems with complex human preferences is a growing priority in fields such as robotics [1, 2], recommendation systems [3], and language models [4, 5, 6]. A key challenge in this endeavor is how to infer and represent such preferences accurately, particularly when they are only available through incomplete signals like pairwise comparisons. This has prompted reinforcement learning from human feedback (RLHF), which has become a widely used framework for preference learning [6, 7]. RLHF streamlines the alignment process by first learning a reward model that assigns scalar scores to different alternatives, typically trained using maximum likelihood estimation under the Bradley–Terry (BT) model. In the second stage, a policy is optimized through reinforcement learning to maximize the expected rewards, guiding the system toward behaviors aligned with human preferences.
Despite its practical success and simplicity, the standard RLHF framework rests on a critical assumption that complex human preferences can be captured by a single scalar reward. Recent research highlights that this assumption often breaks down, especially when human feedback reflects inconsistent or conflicting judgments across evaluators [8]. In particular, RLHF struggles in scenarios involving intransitive or cyclic preferences, where no clear ranking among alternatives can be established, leading to failures in accurately modeling the underlying preferences [9, 10]. To address these limitations, a game-theoretic framework called Nash learning from human feedback (NLHF) has been introduced [9, 10, 11, 12]. NLHF reframes preference learning as a two-player constant-sum game and identifies equilibrium policies that no competing policy can outperform, regardless of the complexity of the underlying preferences.
Nevertheless, the NLHF framework remains limited in its ability to address another critical limitation of single reward model-based alignment approaches: the proportional representation of evaluator preferences. When preferences are aggregated across multiple evaluator groups with distinct viewpoints, both RLHF and NLHF tend to yield policies that do not adequately reflect the full distribution of the evaluator population [8]. To address these challenges, recent research has turned to social choice theory–oriented approaches, such as maximizing the minimum satisfaction across evaluator groups [8, 13] and optimizing social welfare functions [14, 15]. However, these methods generally assume explicit knowledge or clear labels of evaluator groups, which limits their practical applicability since group identities are often implicit or unobservable in real-world. Motivated by this limitation, our research aims to achieve equitable alignment without requiring additional information about the evaluator profile.
Our approach builds upon recent works addressing diverse preference aggregation through an axiomatic approach from social choice theory [16, 17, 18, 19, 20, 12]. Specifically, we propose a novel preference learning algorithm that satisfies two foundational axioms, monotonicity (ensuring that improving an alternative’s ranking cannot decrease its probability) and Pareto efficiency (ensuring that if an alternative is preferred by all, it is favored by the policy), as well as two new axioms we introduce: population-proportional representation (PPR) and population-bounded robustness (PBR). The first new axiom, PPR, ensures the output policy is at least weakly proportional to the true evaluator population shares. This axiom addresses a shortcoming within both the RLHF and NLHF frameworks, namely, their insufficient representation of the population distribution of preferences. The second axiom, PBR, bounds the incentive for manipulation as an affine function of the true population share, thereby guaranteeing robustness. Recent studies have highlighted that conventional preference learning methods are susceptible to strategic misreporting [21]. Unlike existing approaches that incorporate explicit mechanism design to ensure strict strategyproofness [22, 23, 24, 25, 21], our method inherently limits manipulative advantage by constraining policy selection based on estimated feasible population distributions. Further details on related work are provided in Appendix A.
1.1 Our contribution
The first key contribution of this work is demonstrating that the set of feasible population distributions of evaluators can be inferred directly from pairwise comparison data. Leveraging this insight, we develop a novel preference learning framework designed to align policies proportionally with the underlying population distribution. To establish a rigorous theoretical basis, we adopt an axiomatic approach, proving that our framework satisfies two fundamental axioms, monotonicity and Pareto efficiency, and two newly introduced axioms, PPR and PBR. For practical deployment, we present a scalable algorithm with function approximation, allowing our framework to scale to high-dimensional settings such as large language models (LLMs). Finally, the proposed framework is validated through empirical evaluations in both tabular and function approximation settings.
Organization of the paper
In Section 2, we formalize the setting of preference learning and probabilistic social choice, and establish connections between them. In Section 3, motivated by a simple negative example, we introduce two desirable axioms alongside two fundamental axioms. In Section 4, we propose a novel tabular algorithm that satisfies these axioms and provide a theoretical analysis. Section 5 extends our approach to high-dimensional settings using function approximation, enabling scalable offline preference learning. Finally, Section 6 presents empirical evaluations that demonstrate the effectiveness and scalability of our method.
2 Preliminaries
Notation.
We use $[M]:=\{1,2,\dots,M\}$ to denote the set of integers from $1$ to $M$. The set $\Delta(\mathcal{Y})$ represents the probability simplex over a finite set $\mathcal{Y}$. The maximum norm is denoted by $\|\cdot\|_{\infty}$.
2.1 Probabilistic social choice function and preference learning
We begin by reviewing key concepts from social choice theory and preference learning to establish a foundation for our subsequent analysis. Consider a set of $M$ alternatives, denoted by $\mathcal{Y}:=\{y_{1},y_{2},\dots,y_{M}\}$, where each $y\in\mathcal{Y}$ may represent a response generated by a language model or an action in a decision-making task.
We assume each evaluator has a strict and complete ranking over the alternatives, and let $S$ denote the set of all possible rankings (i.e., permutations of $\mathcal{Y}$). Each ranking is represented by $r\in S$, where $r(y_{i})=k$ indicates that alternative $y_{i}$ is ranked $k$-th under $r$. A profile $\sigma\in\Delta(S)$ is a probability distribution over the set of all rankings, where $\sigma_{r}$ represents the proportion of the population that adheres to ranking $r$.
A probabilistic social choice function (PSCF) is a mapping $\Phi:\Delta(S)\to\Delta(\mathcal{Y})$ that assigns to each profile $\sigma$ a policy $\pi$, which is a probability distribution over the alternatives in $\mathcal{Y}$. In practice, however, acquiring a complete profile $\sigma$ is often infeasible due to the high cost of collecting full rankings over a large set of alternatives.
To address this limitation, pairwise preference learning algorithms have been developed, allowing alignment based solely on pairwise comparison data. We define a preference function ${\rm{P}}:\mathcal{Y}^{2}\to[0,1]$, where ${\rm{P}}(y\succ y^{\prime})$ denotes the probability that alternative $y$ is preferred over $y^{\prime}$. Given a profile $\sigma$, let ${\rm{P}}^{\sigma}$ be the preference function induced by the population distribution $\sigma$ over rankings, defined as
${\rm{P}}^{\sigma}(y\succ y^{\prime}):=\sum_{r\in S}\sigma_{r}\cdot\mathbf{1}_{\{r(y)<r(y^{\prime})\}},$ (1)
where $\mathbf{1}_{{r(y)<r(y^{\prime})}}$ is an indicator function equal to $1$ if ranking $r$ places alternative $y$ in a better (i.e., lower) position than $y^{\prime}$, and 00 otherwise. This function captures the expected pairwise preference between $y$ and $y^{\prime}$ under the distribution $\sigma$.
We define $\mathcal{P}$ as the set of all preference functions induced by some profile $\sigma\in\Delta(S)$:
$\mathcal{P}:=\{{\rm{P}}\mid\exists\sigma\in\Delta(S)\text{ s.t. }{\rm{P}}={\rm{P}}^{\sigma}\}.$ (2)
Any ${\rm{P}}\in\mathcal{P}$ satisfies consistency conditions known as the Block-Marschak inequalities [26], including skew-symmetry: ${\rm{P}}(y\succ y^{\prime})+{\rm{P}}(y^{\prime}\succ y)=1\;\forall y,y^{\prime}\in\mathcal{Y}$. A preference learning algorithm is a mapping $\Psi:\mathcal{P}\to\Delta(\mathcal{Y})$ that assigns a policy to each preference function. Throughout this paper, we say that a preference learning algorithm $\Psi$ implements a PSCF $\Phi$ if, for every profile $\sigma\in\Delta(S)$, it holds that $\Psi({\rm{P}}^{\sigma})=\Phi(\sigma)$. The relationships between the profile, preference function, and policy are illustrated in Figure 1.
Figure 1: Illustration of the relationships between the profile, preference function, and policy.
2.2 Two standard preference learning algorithms
Next, we introduce two prominent preference learning algorithms and discuss their connections to established concepts from probabilistic social choice theory.
Reinforcement learning from human feedback (RLHF)
A widely used method in preference modeling is the Bradley–Terry (BT) model, which assigns each alternative $y_{i}$ a scalar reward $r_{i}$. Under the BT model, the preference function follows ${\rm{P}}(y_{i}\succ y_{j})=\exp{(r_{i})}/(\exp{(r_{i})}+\exp{(r_{j})})$. The standard RLHF algorithm first estimates these rewards by maximizing likelihood under the BT model, and subsequently optimizes a policy using reinforcement learning to maximize expected rewards.
Recent research [17] has revealed a close connection between the RLHF algorithm and the maximal Borda rule from social choice theory. Given a profile $\sigma$, the Borda score of an alternative $y$ is defined as $B(y):=\sum_{r\in S}\sigma_{r}\cdot(M-r(y))$, assigning points to alternatives based on their rank in each voter’s preference and summing these points to identify the alternative most broadly favored across all voters. As shown in Appendix B, the ranking induced by optimized rewards under the BT model matches precisely with the ranking induced by Borda scores. Consequently, RLHF without regularization selects the alternative that maximizes Borda scores. We denote this RLHF algorithm by $\Psi_{\mathrm{RLHF}}$ and the corresponding maximal Borda rule by $\Phi_{\mathrm{MB}}$, with $\Psi_{\mathrm{RLHF}}$ implementing $\Phi_{\mathrm{MB}}$. Moreover, direct preference optimization (DPO) [27] extends this approach by incorporating KL divergence regularization with respect to a reference policy into the reward function.
Nash learning from human feedback (NLHF)
As highlighted in recent studies [9, 10, 12], RLHF has limitations in scenarios involving intransitive or cyclic preferences. An alternative $y^{*}$ is called a Condorcet winner if it is preferred by a majority over every other alternative, formally stated as ${\rm{P}}(y^{*}\succ y)>0.5$ for all $y\neq y^{*}$. When aggregating preferences across multiple evaluators, scenarios without a Condorcet winner can arise, which is called the Condorcet paradox. In such cases, selecting the alternative with the highest Borda score fails to adequately represent collective evaluator preferences [9, 10, 12]. To be more specific, in the absence of a clear consensus regarding the socially preferred alternative, the selection of a deterministic policy is incapable of reflecting this lack of agreement (or nuanced preferences).
To address intransitive preferences, maximal lotteries (ML) [28] from probabilistic social choice have been adopted to preference learning. ML formulates preference learning as a two-player constant-sum game $\max_{\pi_{1}\in\Delta(\mathcal{Y})}\min_{\pi_{2}\in\Delta(\mathcal{Y})}{\mathbb{E}}_{(y_{1},y_{2})\sim(\pi_{1},\pi_{2})}\left[{\rm{P}}(y_{1}\succ y_{2})\right]$, where the objective function represents the probability that an alternative drawn from policy $\pi_{1}$ is preferred over one drawn from policy $\pi_{2}$. The equilibrium solution $\pi^{*}$ ensures that no other policy can uniformly outperform it. This game-theoretic approach, termed Nash learning from human feedback (NLHF) [9, 10, 11, 12], is denoted by $\Psi_{\mathrm{NLHF}}$ and implements the ML rule, denoted by $\Phi_{\mathrm{ML}}$.
3 An axiomatic approach to population-proportional preference learning
3.1 Motivating example with binary alternatives
Despite their practical utility, neither RLHF nor NLHF guarantees alignment proportional to the evaluator’s preferences. To illustrate this issue, we present a simple scenario involving binary alternatives.
Consider two alternatives, $\mathcal{Y}=\{y_{1},y_{2}\}$, and a profile $\sigma$ consisting of two distinct groups of evaluators: group $G_{1}$ prefers alternative $y_{1}$ over $y_{2}$, while group $G_{2}$ prefers $y_{2}$ over $y_{1}$. Let $w^{\sigma}_{1}$ and $w^{\sigma}_{2}$ denote the population shares of groups $G_{1}$ and $G_{2}$, respectively.
Suppose the two alternatives are nearly tied, with population distribution given by $(w^{\sigma}_{1},w^{\sigma}_{2})=\left(1/2+\epsilon,1/2-\epsilon\right)$ for an arbitrarily small positive scalar $\epsilon$. Then, the corresponding preference function is given by ${\rm{P}}^{\sigma}(y_{1}\succ y_{2})=1/2+\epsilon$ and ${\rm{P}}^{\sigma}(y_{2}\succ y_{1})=1/2-\epsilon$. Despite this minimal margin $\epsilon$, both algorithms $\Psi_{\mathrm{RLHF}}({\rm{P}}^{\sigma})$ and $\Psi_{\mathrm{NLHF}}({\rm{P}}^{\sigma})$ yield a deterministic policy that select the alternative with slightly greater support, namely $y_{1}$.
This binary example highlights two potential limitations of RLHF and NLHF frameworks. First, selecting policies that focus entirely on a single alternative may not accurately represent preferences across evaluators, raising concerns about bias. Second, these methods have high sensitivity to small perturbations in preference function. Specifically, a slight shift in $\epsilon$ from negative to positive abruptly flips the policy outcome $(\pi(y_{1}),\pi(y_{2}))$ from $(0,1)$ to $(1,0)$, making such approaches vulnerable to small perturbations. These limitations underscore the need for a novel approach that reflects the ratio of $(w_{1}^{\sigma},w_{2}^{\sigma})$ in the resulting policy.
3.2 Axiomatic framework in social choice theory
Social choice theory studies the aggregation of individual preferences, typically employing an axiomatic approach. This involves specifying normatively desirable properties called axioms and precisely characterizing aggregation rules that satisfy these properties. Following this approach, we first introduce the axioms that a PSCF $\Phi$ should satisfy. We then propose a preference learning algorithm $\Psi$ that implements a PSCF meeting all of these axioms.
Fundamental axioms
We begin by reviewing two fundamental axioms in social choice theory: monotonicity and Pareto efficiency. For detailed discussions of these axioms, we refer readers to [29, 20].
Definition 3.1 (Monotonicity).
A PSCF $\Phi$ satisfies monotonicity if, for any alternative $y\in\mathcal{Y}$, improving its ranking in a profile without changing other relative rankings cannot decrease its probability in the resulting policy. Formally, if profile $\sigma^{\prime}$ is obtained from $\sigma$ by improving the ranking of $y$ in some $r\in S$, then $\Phi(\sigma^{\prime})(y)\geq\Phi(\sigma)(y)$.
Definition 3.2 (Pareto efficiency).
A PSCF $\Phi$ satisfies Pareto efficiency if, whenever an alternative $y$ is ranked above $y^{\prime}$ in every ranking $r$ with nonzero population share, the resulting policy assigns at least as much probability to $y$ as to $y^{\prime}$. Formally, if $r(y)<r(y^{\prime})$ for all $r\in S$ with $\sigma_{r}>0$, then $\Phi(\sigma)(y)\geq\Phi(\sigma)(y^{\prime})$.
Proposed axioms
Motivated by the earlier example, we next introduce a new axiom designed to ensure alignment with population distribution of preferences. Let $G_{k}:=\{r\in S\mid r(y_{k})=1\}$ denote the set of rankings in which alternative $y_{k}$ is ranked first. The population share of group $G_{k}$ is denoted by $w^{\sigma}_{k}:=\sum_{r\in G_{k}}\sigma_{r}$. For notational convenience, we define $\sigma_{k}\in\Delta(S)$ as the normalized sub-profile restricted to rankings in $G_{k}$, where $\sigma_{k,r}=\sigma_{r}/w_{k}^{\sigma}$ for all $r\in G_{k}$, and $\sigma_{k,r}=0$ for all $r\notin G_{k}$. Let ${\rm{P}}^{\sigma}_{k}$ denote the group-specific preference function, generated from $\sigma_{k}$, using the mapping defined in (1). By construction, ${\rm{P}}^{\sigma}_{k}(y_{k}\succ y)=1$ for all $y\neq y_{k}$, since this group unanimously prefers $y_{k}$ over all other alternatives. The overall preference function is then a weighted aggregation of the group-specific preferences: ${\rm{P}}^{\sigma}=\sum_{k=1}^{M}w^{\sigma}_{k}{\rm{P}}^{\sigma}_{k}$.
Under this definition, our first axiom ensures that the resulting policy proportionally reflects the underlying population shares of each group.
Definition 3.3 ($\alpha$-Population-proportional representation ($\alpha$-PPR)).
A PSCF $\Phi$ satisfies $\alpha$-population-proportional representation if $\pi(y_{k})/w^{\sigma}_{k}\geq\alpha(\sigma)$ for all $\sigma\in\Delta(S)$ and $y_{k}\in\mathcal{Y}$, where $\pi=\Phi(\sigma)$ and $\alpha:\Delta(S)\to(0,1]$.
The function $\alpha(\sigma)$ quantifies the strength of alignment: a higher value of $\alpha$ implies stronger alignment with $w^{\sigma}$, with $\alpha(\sigma)=1$ indicating perfect proportional representation.
While the primary focus of this paper is the $\alpha$-PPR axiom, we also examine the robustness of $\Phi$ against manipulation through the following axiom.
Definition 3.4 (Single-group manipulated profile).
Given a profile $\sigma$ and a group index $k\in[M]$, a profile $\sigma^{\prime}_{k}$ is called a single-group manipulated profile of $\sigma$ if $\sigma^{\prime}_{k}$ can be obtained by modifying only the ranking distribution of the sub-profile $\sigma_{k}$. Formally, $\sigma^{\prime}_{k}$ is a single-group manipulated profile of $\sigma$ if there exists a profile $\sigma^{\prime}$ such that $\sigma^{\prime}_{k}=\sigma+w_{k}^{\sigma}(\sigma^{\prime}-\sigma_{k})$.
Definition 3.5 ($\gamma$-Population-bounded robustness ($\gamma$-PBR)).
A PSCF $\Phi$ satisfies $\gamma$-population-bounded robustness if, for any profile $\sigma$ and its single-group manipulated profile $\sigma^{\prime}_{k}$, we have $\Phi(\sigma^{\prime}_{k})(y_{k})\leq\gamma_{1}w_{k}^{\sigma}+\gamma_{2}$, where $\gamma=(\gamma_{1},\gamma_{2})$, $\gamma_{1}>0$, and $\gamma_{1}+\gamma_{2}=1$.
The $\gamma$-PBR axiom ensures that the maximum influence a single group can exert through manipulation is bounded above by an affine function of its population share. Specifically, a group can only achieve a deterministic policy selection for its preferred alternative (i.e., $\Phi(\sigma^{\prime}_{k})(y_{k})=1$) only if it constitutes the entire evaluator population (i.e., $w_{k}^{\sigma}=1$). Note that a larger $\gamma_{1}$ provides a stronger robustness guarantee. Particularly, $\gamma_{1}=1$ implies that the manipulated policy value is limited exactly to the group’s true population share.
It is also important to emphasize that the proposed $\gamma$-PBR axiom is not equivalent to strategyproofness as defined in social choice theory. Specifically, $\gamma$-PBR considers manipulation by a single group while assuming fixed profiles for other groups, whereas strategyproofness assumes each group can strategically manipulate its own sub-profile independently.
We next demonstrate that the standard PSCFs either fail to satisfy the proposed axioms or are not implementable by a preference learning algorithm. Consider a PSCF that aligns the policy exactly with each group’s population distribution, commonly referred to as a random dictatorship in probabilistic social choice theory [29].
Definition 3.6 (Random dictatorship).
A PSCF $\Phi_{\mathrm{RD}}$ is called a random dictatorship if $\Phi_{\mathrm{RD}}(\sigma)=w^{\sigma}$ for all $\sigma\in\Delta(S)$.
By definition, $\Phi_{\mathrm{RD}}$ satisfies both proposed axioms in their strongest forms: $\alpha$-PPR with $\alpha(\sigma)=1$ for all $\sigma\in\Delta(S)$, and $\gamma$-PBR with $\gamma=(1,0)$.
The following proposition establishes that $\Phi_{\mathrm{MB}}$ and $\Phi_{\mathrm{ML}}$ violate even the weakest forms of these axioms, whereas $\Phi_{\mathrm{RD}}$ satisfies all four axioms. The proof is provided in Appendix C.
Proposition 3.7.
$\Phi_{\mathrm{MB}}$ and $\Phi_{\mathrm{ML}}$ violate the $\alpha$-PPR axiom for any $\alpha$ and the $\gamma$-PBR axiom for any $\gamma$. $\Phi_{\mathrm{RD}}$ satisfies all four axioms.
Finally, we note that $\Phi_{\mathrm{RD}}$ cannot be implemented by a pairwise preference learning algorithm. This is because there exist distinct profiles $\sigma_{1}$ and $\sigma_{2}$ that result in identical preference functions ${\rm{P}}^{\sigma_{1}}={\rm{P}}^{\sigma_{2}}$ but different population shares $w^{\sigma_{1}}\neq w^{\sigma_{2}}$. A proof with an illustrative example is provided in Appendix D. Thus, since the exact value of $w^{\sigma}$ cannot be determined solely from observing ${\rm{P}}^{\sigma}$, $\Phi_{\mathrm{RD}}$ cannot be implemented through a mapping from preference functions to policies. Consequently, our goal is to construct a preference learning algorithm $\Psi$ that implements a PSCF $\Phi$ satisfying all four axioms.
4 Algorithmic framework and theoretical guarantees
4.1 Population distribution recovery from pairwise preferences
In this section, we introduce a preference algorithm $\Psi$, which implements a PSCF satisfying all four axioms presented in the previous section. The framework first estimates the feasible set of underlying population distributions $w$ from given pairwise preferences ${\rm{P}}$, and subsequently constructs a policy $\pi$ closely aligned with the inferred $w$. We begin with the definition of a feasible population distribution and the characterization of the set of all feasible population distributions.
Definition 4.1.
A population distribution $w$ is considered feasible given ${\rm{P}}$, if there exists a profile $\sigma\in\Delta(S)$ such that $w=w^{\sigma}$ and ${\rm{P}}={\rm{P}}^{\sigma}$.
Proposition 4.2.
The set of all feasible population distributions given ${\rm{P}}$ can be expressed as
$\begin{split}\mathcal{W}({\rm{P}}):=\Big{\{}w\in\Delta(\mathcal{Y})\;\Big{|}\;\exists({\rm{P}}_{1},\ldots,{\rm{P}}_{M})\in\mathcal{P}^{M}\;\text{s.t.}\;&{\rm{P}}=\sum_{i=1}^{M}w_{i}{\rm{P}}_{i},\\ &{\rm{P}}_{i}(y_{i}\succ y)=1\;\forall y\in\mathcal{Y},i\in[M]\Big{\}}.\end{split}$ (3)
See Appendix E for the proof. In words, a population distribution $w$ is feasible if and only if there exists a collection of group-specific preference function $({\rm{P}}_{1},\ldots,{\rm{P}}_{M})$ such that ${\rm{P}}$ is their weighted aggregation, and each ${\rm{P}}_{i}$ reflects a group of evaluators who unanimously prefer $y_{i}$ over all other alternatives.
The exact characterization of the set $\mathcal{W}({\rm{P}})$ is challenging due to the constraints imposed by the set $\mathcal{P}$. To address this, we propose a tractable polyhedral outer approximation of the set $\mathcal{W}({\rm{P}})$, with the number of constraints growing only linearly with the dimension $M$.
Definition 4.3.
For each $i\in[M]$, define $u_{i}:=\min_{y\in\mathcal{Y}\setminus\{y_{i}\}}{\rm{P}}(y_{i}\succ y)$.
Theorem 4.4.
The set of all feasible population distributions satisfies
$\mathcal{W}({\rm{P}})\subseteq\overline{\mathcal{W}}({\rm{P}}):=\left\{w\in\Delta(\mathcal{Y})\;\Big{|}\;w_{i}\leq u_{i}\;\forall i\in[M]\right\}.$ (4)
The proof is provided in Appendix F. To provide intuition for this theorem: $u_{i}$ can be rewritten as $1-\max_{y\neq y_{i}}{\rm{P}}(y\succ y_{i})=1-{\rm{P}}(y^{*}\succ y_{i})$, where $y^{*}$ denotes the alternative that is preferred over $y_{i}$ by the largest proportion of evaluators. Then, $u_{i}$ represents the remaining population share after excluding those who prefer $y^{*}$ to $y_{i}$. Thus, $w_{i}$ cannot exceed this value, as the $w_{i}$ proportion of evaluators would always report $y_{i}$ as their preferred option. We note that $\overline{\mathcal{W}}({\rm{P}})$ provides a tight approximation (i.e., $\mathcal{W}({\rm{P}})=\overline{\mathcal{W}}({\rm{P}})$) if we consider an extended profile setting (see Appendix F for details).
Proposition 4.2 and Theorem 4.4 indicate a fundamental limitation in preference learning: the true value $w^{\sigma}$ is not uniquely identifiable from pairwise comparison data in general, unless the set $\mathcal{W}({\rm{P}})$ contains only a single element. As a result, achieving perfect population-proportional representation (i.e., $\alpha(\sigma)=1$) is generally impossible. This motivates the design of preference learning algorithms that aim to satisfy $\alpha$-PPR for the largest achievable value of $\alpha$.
4.2 Proposed algorithmic framework with axiomatic guarantees
Given a polyhedron $\overline{\mathcal{W}}({\rm{P}})$, our goal is to select a policy $\pi$ that guarantees the proportional alignment to all $w\in\overline{\mathcal{W}}({\rm{P}})$. To this end, we propose to assign probabilities to alternatives in proportion to the derived upper bounds $u_{i}$.
Definition 4.5.
The preference learning algorithm $\Psi^{*}$ maps a preference function ${\rm{P}}$ to the following policy:
$\pi(y_{i})=\frac{u_{i}}{\sum_{j=1}^{M}u_{j}},\quad\forall i\in[M].$ (5)
Let $\Phi^{*}$ denote the PSCF implemented by $\Psi^{*}$.
This construction adopts a conservative strategy for handling uncertainty in $w^{\sigma}$ by assigning probabilities proportional to the most conservative estimate of each $w^{\sigma}_{i}$. By doing so, the algorithm minimizes the worst-case misalignment caused by the inevitable information loss from pairwise comparisons. Formally, it solves $\max_{\pi\in\Delta(\mathcal{Y})}\min_{w\in\overline{\mathcal{W}}({\rm{P}})}\|\pi/w\|_{\infty}$.
We first establish the foundational axiomatic guarantees of the proposed framework. The proofs are provided in Appendix H.
Theorem 4.6 (Monotonicity & Pareto efficiency).
The proposed PSCF $\Phi^{*}$ satisfies the monotonicity and the Pareto efficiency.
Next, we show that $\Phi^{*}$ satisfies the $\alpha$-PPR axiom. The following lemma establishes that the ratio between the resulting policy and the true population share is lower bounded by the inverse of the total sum of the upper bounds $u_{i}$.
Lemma 4.7.
For any profile $\sigma\in\Delta(S)$, the policy $\pi=\Phi^{*}(\sigma)$ satisfies
$\frac{\pi(y_{k})}{w_{k}^{\sigma}}\geq\left(\sum_{i=1}^{M}u_{i}\right)^{-1}\forall k\in[M].$ (6)
The next lemma shows that this lower bound is always finite and depends on the number of non-dominated alternatives in the profile.
Definition 4.8 ($\delta$-dominated alternative).
For any $\delta\in[0,1]$, an alternative $y\in\mathcal{Y}$ is said to be $\delta$-dominated in a profile $\sigma$ if there exists an alternative $y^{\prime}\in\mathcal{Y}\setminus\{y\}$ such that ${\rm{P}}^{\sigma}(y^{\prime}\succ y)\geq\delta$.
Lemma 4.9.
Let $w^{\sigma,1}$ and $w^{\sigma,2}$ denote the largest and second-largest elements of $w^{\sigma}$, respectively. Consider any $\delta\in[0,1]$, and let $N^{\sigma}_{\delta}$ be the number of alternatives that are not $\delta$-dominated in profile $\sigma$. Then the lower bound in (6) lies within the range $[\alpha(\sigma),1]$, where
$\alpha(\sigma):=\left[(N^{\sigma}_{\delta}-1)(1-w^{\sigma,1})+(1-w^{\sigma,2})+(M-N^{\sigma}_{\delta})(1-\delta)\right]^{-1}.$ (7)
See Appendix I for the proofs of Lemma 4.7 and Lemma 4.9. Combining both Lemmas, we obtain the following $\alpha$-PPR guarantee:
Theorem 4.10 ($\alpha$-PPR guarantee).
The PSCF $\Phi^{*}$ satisfies the $\alpha$-PPR axiom with $\alpha$ defined in (7).
Lemma 4.7 suggests that the actual alignment performance improves as $\sum_{j=1}^{M}u_{j}$ approaches 1. This typically occurs when the number of non-$1$-dominated alternatives is small. Notably, when there are only two non-$1$-dominated alternatives, substituting $N^{\sigma}_{1}=2$ and $w^{\sigma,1}+w^{\sigma,2}=1$ into Equation (7) yields $\alpha(\sigma)=1$. This implies that the proposed algorithm can recover the exact population distribution in such cases. Moreover, when the population is dominated by a single dominating group, meaning $(w^{\sigma,1},w^{\sigma,2})$ approaches $(1,0)$, then $\alpha(\sigma)$ also approaches 1. Importantly, because $\sum_{j=1}^{M}u_{j}$ can be computed directly from a given preference function ${\rm{P}}$, the alignment accuracy of the resulting policy can be evaluated at test time.
Finally, we present the population-bounded robustness of the proposed method.
Theorem 4.11 ($\gamma$-PBR guarantee).
Let $\pi^{\prime}=\Phi^{*}(\sigma^{\prime}_{k})$ denote a policy resulting from single-group manipulation by group $G_{k}$. Then, the following inequality holds:
$\pi^{\prime}(y_{k})\leq\frac{u_{k}}{u_{k}+1-w^{\sigma}_{k}}\leq\frac{1}{2}(w_{k}^{\sigma}+1).$ (8)
Thus, the PSCF $\Phi^{*}$ satisfies $\gamma$-PBR with $(\gamma_{1},\gamma_{2})=(1/2,1/2)$.
The proof is provided in Appendix J. Note that $(\gamma_{1},\gamma_{2})=(1/2,1/2)$ represents a worst-case bound. The actual manipulability for each group is more tightly bounded by $u_{k}/(u_{k}+1-w_{k}^{\sigma})$. For instance, if $u_{k}\leq 1/2$ and $w_{k}^{\sigma}\leq 1/2$, then $\pi^{\prime}(y_{k})\leq 1/2$. This indicates that a non-majority group cannot elevate their preferred alternative to majority status through manipulation.
As discussed previously, $\gamma$-PBR differs from strategyproofness in its definition. However, the above result can be extended to provide a weaker form of strategyproofness (see Appendix K for further details).
4.3 Balancing population-proportional representation and Condorcet consistency
One potential limitation of the proposed framework is that $\Phi^{*}$ does not satisfy Condorcet consistency, a fundamental axiom which requires always selecting a Condorcet winner if one exists.
Definition 4.12 (Condorcet consistency).
A PSCF $\Phi$ satisfies Condorcet consistency if, for any profile $\sigma$ with a Condorcet winner $y^{*}$, $\Phi(\sigma)(y^{*})=1$.
In fact, it is impossible for any method to simultaneously satisfy both $\alpha$-PPR and Condorcet consistency (see Appendix L for the proof).
Proposition 4.13.
No PSCF can simultaneously satisfy $\alpha$-PPR and Condorcet consistency.
To balance two axioms, we define a softmax-relaxed algorithm $\Psi^{\beta}$ (and its corresponding PSCF $\Phi^{\beta}$), by modifying $\Psi^{*}$ as follows:
$\pi(y_{i})=\frac{u_{i}\exp(\beta u_{i})}{\sum_{j=1}^{M}u_{j}\exp(\beta u_{j})}.$ (9)
The parameter $\beta\geq 0$ controls how sharply the policy concentrates on alternatives with higher $u_{i}$ values. When $\beta=0$, the algorithm reduces to the original $\Psi^{*}$. As $\beta\to\infty$, the policy becomes deterministic and converges to $\pi(y^{*})=1$, where $y^{*}=\operatorname*{arg\,max}_{i\in[M]}u_{i}$. This limiting $\Phi^{\infty}$ can be interpreted as selecting the alternative that is most robust to adversarial alternative selection, and it satisfies Condorcet consistency (see Appendix M for the proof).
Proposition 4.14.
$\Phi^{\infty}$ satisfies Condorcet consistency.
The softmax relaxation offers a smooth trade-off between $\alpha$-PPR and Condorcet consistency, mediated by the parameter $\beta$. We analyze the behavior of intermediate $\beta$ values in Appendix N, and empirically demonstrate the effects of varying $\beta$ in Section 6.
4.4 Connection to pairwise majority consistency (PMC)
Finally, we discuss the connection to pairwise majority consistency (PMC) [20], which imposes a stronger consistency requirement, ensuring the entire policy ranking aligns with majority preferences.
Definition 4.15 (Pairwise majority consistent ranking (PMC ranking)).
A ranking $r^{\sigma}$ is a called a PMC ranking of a profile $\sigma$ if for all $y_{i},y_{j}\in\mathcal{Y}$, a majority of evaluators prefer alternative $y_{i}$ to alternative $y_{j}$ in $\sigma$ if and only if $y_{i}$ is ranked higher than $y_{j}$ in $r^{\sigma}$. Formally, ${\rm{P}}^{\sigma}(y_{i}\succ y_{j})>1/2$ if and only if $r^{\sigma}(y_{i})<r^{\sigma}(y_{j})$.
Definition 4.16 (Pairwise majority consistency (PMC)).
A PSCF $\Phi$ satisfies PMC if, for any profile $\sigma\in\Delta(S)$ that has a PMC ranking $r^{\sigma}$, $\Phi(\sigma)$ has the same ranking with $r^{\sigma}$, i.e. $\Phi(\sigma)(y_{i})\geq\Phi(\sigma)(y_{j})$ if $r^{\sigma}(y_{i})<r^{\sigma}(y_{j})$.
It can be shown that any $\Phi^{\beta}$ with finite $\beta\geq 0$ violates PMC, and only the limiting PSCF $\Phi^{\beta}$ satisfies PMC (see Appendix O for the proof).
Proposition 4.17.
Any $\Phi^{\beta}$ with finite $\beta\geq 0$ violates PMC. $\Phi^{\infty}$ satisfies PMC.
While $\Phi^{\beta}$ approximately satisfies PMC as $\beta\to\infty$ if we allow some slack in the rankings (e.g., $\Phi(\sigma)(y_{i})\geq\Phi(\sigma)(y_{j})-\epsilon$ for some small $\epsilon>0$) in the definition of PMC. However, exploring this approximate consistency is beyond the scope of this paper and is left for future research.
5 Scalable offline algorithm with function approximation
5.1 Offline pairwise comparison dataset
In practical applications of preference learning, the preference function often depends on additional context or state, denoted by $x\in\mathcal{X}$. For instance, in LLMs, $x$ represents the input prompt or conversational history that provides the specific context for generating a preferred response. Accordingly, we define the context-dependent preference function as ${\rm{P}}(\cdot\succ\cdot\mid\cdot):\mathcal{Y}^{2}\times\mathcal{X}\to[0,1]$, which is unknown and must be estimated from empirical data. We consider an offline dataset of pairwise comparisons $\mathcal{D}=\{(x_{i},y_{i}^{w},y_{i}^{l})\}_{i=1}^{N}$, where $y_{i}^{w}$ is preferred over $y_{i}^{l}$ under context $x_{i}$. Each query is assumed to be drawn i.i.d. from a joint distribution of $\rho(x)$ and $\pi_{d}(y\mid x)$, and labeled according to the preference function ${\rm{P}}$. Our goal is to use this offline dataset to learn a policy $\pi:\mathcal{X}\mapsto\Delta(\mathcal{Y})$ following the framework introduced in the previous sections.
5.2 Two-Phase offline preference learning algorithm
To be specific, we aim to approximate a softmax policy proposed in Section 4.3:
$\pi(y\mid x):=\frac{u(y\mid x)\exp(\beta u(y\mid x))}{\sum_{y\in\mathcal{Y}}u(y\mid x)\exp(\beta u(y\mid x))},\quad\text{where}\quad u(y\mid x):=\min_{z\in\mathcal{Y}}\;{\rm{P}}(y\succ z\mid x).$ (10)
We propose a two-phase algorithm that first estimates $u$ and then estimates $\pi$ based on $u$.
Phase 1: Estimating $u$.
To estimate $u$, we first train the $\mu$ model using the following loss function, the offline dataset $\mathcal{D}$, and the parameterized function class $\mathcal{F}_{\mu}$:
$\hat{\mu}\in\operatorname*{arg\,min}_{\mu\in{\mathcal{F}}_{\mu}}\frac{1}{N}\sum_{i=1}^{N}\frac{\mu(y^{l}_{i}\mid x_{i},y^{w}_{i})}{\pi_{d}(y^{l}_{i}\mid x_{i})}.$ (11)
Then, the estimated $\hat{u}$ can be obtained from $\hat{u}(y\mid x)=\sum_{z\in\mathcal{Y}}\hat{{\rm{P}}}(y\succ z\mid x)\hat{\mu}(z\mid x,y)$, where $\hat{{\rm{P}}}$ denotes the empirical estimate of the preference function (see Appendix P for derivation).
Phase 2: Estimating $\pi$.
Let $\hat{\pi}_{\beta}$ be the normalized softmax policy constructed with $\hat{u}$ following (10). In the second phase, the policy model is trained by minimizing the distance to $\hat{\pi}_{\beta}$ over a function class $\mathcal{F}_{\pi}$:
$\hat{\pi}\in\operatorname*{arg\,min}_{\pi\in\mathcal{F}_{\pi}}\mathbb{E}_{x\sim\rho}\Big{[}L^{\pi}(\pi(\cdot\mid x),\hat{\pi}_{\beta}(\cdot\mid x))\Big{]}.$ (12)
Here, $L^{\pi}$ denotes a divergence or distance metric between two policies. An example using KL divergence, along with other practical considerations for LLM applications, is provided in Appendix Q.
6 Experiments
6.1 Tabular experiment: movie recommendation
Datasets and experimental setup.
To verify our theoretical findings in the tabular setting, we consider a movie recommendation task. The profile, consisting of $|\sigma|=1,297$ human evaluators’ rankings over $M=20$ movies, is extracted from the MovieLens 1M dataset [30]. In each episode, $N=10^{5}$ pairwise comparisons are sampled uniformly, and a policy is trained using the proposed algorithm $\Psi^{\beta}$, along with two baselines: $\Psi_{\mathrm{RLHF}}$ and $\Psi_{\mathrm{NLHF}}$. We evaluate two alignment metrics, averaged over $50$ episodes: (i) the win rate against the uniform policy, and (ii) the PPR level, $\alpha(\sigma)=\min_{i\in[M]}(\pi(y_{i})/w^{\sigma}_{i})$ (where a higher $\alpha(\sigma)$ indicates better PPR). To evaluate PBR, we allow a single evaluator group to manipulate its rankings to maximize its policy gain, $\pi^{\prime}(y_{k})-\pi(y_{k})$, and average it across all groups.
Results and discussion.
The left panel of Figure 2 shows the average and standard deviation of the win rate and PPR level across different preference learning algorithms. As anticipated, both RLHF and NLHF achieve high win rates of $0.7784$ and $0.7768$, respectively, but both exhibit a PPR level of 00. For our proposed algorithm $\Psi^{\beta}$, we observed the predicted trade-off: as $\beta$ increases, the win rate improves from $0.5987$ to $0.7784$, while the PPR level correspondingly decreases from $0.4869$ to 00. The average value of $u_{i}$ was $0.1892$, suggesting that the set $\overline{\mathcal{W}}({\rm{P}})$ (4) provides a meaningful estimation of $w^{\sigma}$ in our method. Regarding PBR, the average gain was calculated as $0.0552$ for RLHF, $0.0166$ for NLHF, and $8.896\times 10^{-4}$ when $\beta=10^{0}$. Generally, $\Psi^{\beta}$ outperformed the two baseline methods when $\beta\leq 10^{1}$. These findings indicate that our proposed algorithm significantly reduces susceptibility to manipulation, supporting its robustness guarantee.
Figure 2: Tabular experiment results (Section 6.1) for three methods: $\Psi^{\beta}$ (solid line), $\Psi_{\mathrm{RLHF}}$ (dashed horizontal line), and $\Psi_{\mathrm{NLHF}}$ (dotted horizontal line). Left: win rate (left axis, blue) and PPR level (right axis, orange). Right: policy gain through manipulation.
6.2 Large-scale experiment: instruction-tuned LLMs
Datasets and experimental setup.
We next evaluate the scalable algorithm proposed in Section 5, by fine-tuning the Qwen2-1.5B-Instruct model [31] using the Alpaca-GPT4 dataset [32]. Following prior work [33, 8], we divide the evaluators into two groups across three categories and generate pairwise comparison data using the annotation model Llama-3.1-8B-Instruct [34]. Details on the prompts used for data generation and other hyperparameter settings are provided in Appendix R. We evaluate two metrics: (i) the win rate against a reference policy and (ii) the PPR level $\alpha(\sigma)$, and compare them with the baseline method DPO [27]. All experimental code is provided in the supplemental material.
Results and discussions.
Table 1 presents the win rate against the reference policy and the evaluated PPR level $\alpha$ for various algorithms. The results indicate that our algorithm achieves a win rate and PPR level comparable to DPO across different group categories. Notably, we observed that the trade-off between win rate and PPR level within our proposed algorithm can be effectively controlled by adjusting the hyperparameter $\beta$, regardless of the category. These results provide encouraging initial evidence for the scalability of our proposed framework in balancing win rate and population-proportional representation.
However, for practical deployment, future research should include more extensive evaluations across diverse datasets and focus on developing more efficient algorithms, as the current two-phase training algorithm incurs higher computational costs compared to DPO. Additionally, in the LLM experiments, the PPR level $\alpha$ cannot be directly observed and must be estimated from the classification results of an annotation model, thus suffering from inherent biases and variances in that model. Developing more accurate and efficient methods for measuring the true PPR level in LLMs is an important and promising direction for future research.
Table 1: Win rate and PPR level ($\alpha$) by algorithm.
Category Metric $\beta=0$ $\beta=10^{-2}$ $\beta=10^{0}$ $\beta=10^{2}$ DPO
Expertise Win rate 0.7613 0.7634 0.7636 0.7664 0.7697
PPR ($\alpha$) 0.5385 0.5204 0.5208 0.5180 0.5176
Informativeness Win rate 0.8266 0.8364 0.8632 0.8692 0.8205
PPR ($\alpha$) 0.1428 0.1392 0.1273 0.1276 0.1421
Style Win rate 0.8443 0.8660 0.8570 0.8718 0.8319
PPR ($\alpha$) 0.7515 0.7445 0.6928 0.7014 0.8014
7 Conclusion and future directions
This paper introduces a novel preference learning framework that aligns policies proportional to population distributions inferred from pairwise comparison data. We believe this framework offers a new perspective on evaluating alignment algorithms, shifting the focus beyond the conventional emphasis on win rate. Furthermore, our work strengthens the connection between preference learning and social choice theory by implementing a new class of probabilistic social choice functions, unlike traditional approaches like the Borda rule. Future research will focus on extending the algorithm to incorporate lower rankings and developing more efficient single-phase algorithms for high-dimensional environments.
8 Acknowledgement
The authors acknowledge the MIT Office of Research Computing and Data for providing high performance computing resources that have contributed to the research results reported within this paper.
Appendix A Additional related work
In this section, we provide an additonal discussion of recent literature that bridges social choice theory and preference learning algorithms designed to capture complex human preferences.
Limitations of the BT model in representing complex human preferences
Recent studies have highlighted significant limitations of the standard RLHF approach under BT model assumption, which often fails to capture the multifaceted and sometimes conflicting nature of human preferences. For example, Kim et al. [35] demonstrated that the standard MLE algorithms under the BT model can become unstable, particularly in the presence of evaluators exhibiting greedy behavior. They proposed to address this limitation by estimating a set of feasible reward functions without relying on specific modeling assumptions. Additionally, Siththaranjan et al. [17] established a theoretical equivalence between RLHF and the Borda voting rule, showing that the optimized rankings from standard methods frequently violate majority preferences. To address this issue, they introduced a distributional approach incorporating hidden context variables to address diverse evaluator preferences. Furthermore, Ge et al. [20] analyzed reward optimization methods under parameterizations, revealing their inherent violation of fundamental axioms such as Pareto efficiency. They proposed a novel algorithm explicitly designed to satisfy these axioms.
Unbiased aggregation of heterogeneous preferences
Parallel research efforts have explored unbiased aggregation of heterogeneous human preferences. Chakraborty et al. [8] formally proved the impossibility of equitably aligning single-reward models across diverse evaluator groups, and proposed learning reward mixtures using the EM algorithm followed by maximizing the minimum utility across subpopulations. Additionally, Zhong et al. [14] conducted a rigorous analysis of multi-group reward learning under various social welfare criteria, such as Nash, utilitarian, and Leximin functions, and provided theoretical alignment guarantees. Park et al. [22] proposed a probabilistic opinion pooling function that directly aggregates multiple probabilistic models into a single policy, as well as personalized algorithms that output individualized policies after estimating confidence sets. While these approaches show promise, they generally rely on explicit group identification, restricting their applicability in scenarios where group labels are unavailable or difficult to determine. In contrast, our work does not require explicit knowledge of evaluator groups. Instead, we infer population distributions directly from pairwise comparison data and align policies accordingly.
Appendix B Equivalence of BT-MLE rewards ranking and Borda ranking
Proposition B.1.
Let $r^{*}\in\mathbb{R}^{M}$ be a maximizer of the likelihood function
$L(r):=\sum_{i<j}\left[{\rm{P}}^{\sigma}(y_{i}\succ y_{j})\log\left(\frac{e^{r_{i}}}{e^{r_{i}}+e^{r_{j}}}\right)+{\rm{P}}^{\sigma}(y_{j}\succ y_{i})\log\left(\frac{e^{r_{j}}}{\,e^{r_{i}}+e^{r_{j}}}\right)\right].$ (13)
Then, the ordering of alternatives induced by $r^{*}$ is identical to the ordering induced by the Borda score $B$ of $\sigma$. Formally, for any $i,j\in[M]$,
$r^{*}_{i}>r^{*}_{j}\quad\Longleftrightarrow\quad B(y_{i})>B(y_{j}).$ (14)
Proof.
The gradient of $L(r)$ with respect to $r_{i}$ is given by:
$\frac{\partial L(r)}{\partial r_{i}}=\sum_{j\neq i}\left[{\rm{P}}^{\sigma}(y_{i}\succ y_{j})-\mathrm{sigmoid}(r_{i}-r_{j})\right],$ (15)
where $\mathrm{sigmoid}(x):=1/(1+e^{-x})$. At the optimal solution $r^{*}$, the first-order condition requires that
$\sum_{j\neq i}\left[{\rm{P}}^{\sigma}(y_{i}\succ y_{j})-\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{j})\right]=0.$ (16)
Now, consider two distinct alternatives $i$ and $k$, and suppose that $r^{*}_{i}>r^{*}_{k}$. Since the sigmoid function is monotonically increasing, for any $j\neq i,k$, we have $\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{j})>\mathrm{sigmoid}(r^{*}_{k}-r^{*}_{j})$, and also $\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{k})>\mathrm{sigmoid}(r^{*}_{k}-r^{*}_{i})$. From the first-order conditions at optimality, we have:
$\sum_{j\neq i}{\rm{P}}^{\sigma}(y_{i}\succ y_{j})=\sum_{j\neq i}\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{j})\;\text{ and }\;\sum_{j\neq k}{\rm{P}}^{\sigma}(y_{k}\succ y_{j})=\sum_{j\neq k}\mathrm{sigmoid}(r^{*}_{k}-r^{*}_{j}).$ (17)
Since $r^{*}_{i}>r^{*}_{k}$, it follows that
$\sum_{j\neq i}\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{j})>\sum_{j\neq k}\mathrm{sigmoid}(r^{*}_{k}-r^{*}_{j}).$ (18)
Therefore, we have
$\sum_{j\neq i}{\rm{P}}^{\sigma}(y_{i}\succ y_{j})>\sum_{j\neq k}{\rm{P}}^{\sigma}(y_{k}\succ y_{j}).$ (19)
By definition, ${\rm{P}}^{\sigma}(y_{i}\succ y_{j})=\sum_{r\in S}\sigma_{r}\cdot\mathbf{1}_{\{r(y_{i})<r(y_{j})\}}$. Substituting this into the inequality above, we get
$\sum_{r\in S}\sigma_{r}\cdot\sum_{j\neq i}\mathbf{1}_{\{r(y_{i})<r(y_{j})\}}>\sum_{r\in S}\sigma_{r}\cdot\sum_{j\neq k}\mathbf{1}_{\{r(y_{k})<r(y_{j})\}}.$ (20)
Recall that the Borda score is defined as $B(y):=\sum_{r\in S}\sigma_{r}\cdot(M-r(y))$, we can rewrite the inner sums in the inequality as:
$\sum_{j\neq i}\mathbf{1}_{\{r(y_{i})<r(y_{j})\}}=(M-1)-(r(y_{i})-1)=M-r(y_{i}),$ (21)
and similarly,
$\sum_{j\neq k}\mathbf{1}_{\{r(y_{k})<r(y_{j})\}}=M-r(y_{k}).$ (22)
Thus, the inequality becomes
$\sum_{r\in S}\sigma_{r}\cdot(M-r(y_{i}))>\sum_{r\in S}\sigma_{r}\cdot(M-r(y_{k})),$ (23)
which is equivalent to $B(y_{i})>B(y_{k})$.
For the converse, assume $B(y_{i})>B(y_{k})$. Following similar steps in reverse, this implies
$\sum_{j\neq i}{\rm{P}}^{\sigma}(y_{i}\succ y_{j})>\sum_{j\neq k}{\rm{P}}^{\sigma}(y_{k}\succ y_{j}),$ (24)
which leads to
$\sum_{j\neq i}\mathrm{sigmoid}(r^{*}_{i}-r^{*}_{j})>\sum_{j\neq k}\mathrm{sigmoid}(r^{*}_{k}-r^{*}_{j}).$ (25)
This inequality can only hold if $r^{*}_{i}>r^{*}_{k}$. Therefore, we have shown that $r^{*}_{i}>r^{*}_{j}\iff B(y_{i})>B(y_{j})$, completing the proof. ∎
Appendix C Proof of Proposition 3.7
We first demonstrate that $\Phi_{\mathrm{MB}}$ and $\Phi_{\mathrm{ML}}$ violate $\alpha$-PPR. Consider a preference profile $\sigma$ with the following characteristics: (i) The population share of each group is nearly identical, with $G_{1}$ having a population share $w^{\sigma}_{1}$ that is $\epsilon$ greater than the average, and $G_{2}$ having a population share $w^{\sigma}_{2}$ that is $\epsilon$ less than the average. (ii) Within each group $G_{k}$, there is indifference between any two alternatives other than $y_{k}$. That is, ${\rm{P}}^{\sigma}_{k}(y_{i}\succ y_{j})=1/2$ for all $i,j\neq k$. Given this profile, we will show that for any $\epsilon>0$, both RLHF and NLHF yield a deterministic policy that selects $y_{1}$.
The population distribution and pairwise preference function satisfy
$w^{\sigma}=\left(\frac{1}{M}+\epsilon,\frac{1}{M}-\epsilon,\frac{1}{M},\frac{1}{M},\dots,\frac{1}{M}\right)\;\text{and}\;{\rm{P}}^{\sigma}_{k}(y_{i}\succ y_{j})=\frac{1}{2},\quad\forall i,j\neq k.$ (26)
Then, the aggregated pairwise preferences ${\rm{P}}^{\sigma}$ are computed as follows:
  • ${\rm{P}}^{\sigma}(y_{1}\succ y_{2})=\frac{1}{2}+\epsilon$
  • ${\rm{P}}^{\sigma}(y_{1}\succ y)=\frac{1}{2}+\frac{\epsilon}{2}$ for any $y\neq y_{1},y_{2}$
  • ${\rm{P}}^{\sigma}(y_{2}\succ y)=\frac{1}{2}-\frac{\epsilon}{2}$ for any $y\neq y_{1},y_{2}$
  • ${\rm{P}}^{\sigma}(y\succ y^{\prime})=\frac{1}{2}$ for any $y,y^{\prime}\neq y_{1},y_{2}$
Under this profile, for any $\epsilon>0$, both $\Phi_{\mathrm{MB}}$ and $\Phi_{\mathrm{ML}}$ result in a policy where $\pi(y_{1})=1$, and $\pi(y_{i})=0$ for any $i\neq 1$. This implies that $\pi(y_{i})/w_{i}^{\sigma}=0$ for any $i\neq 1$, which violates $\alpha$-PPR for any $\alpha>0$.
Next, we show that $\Phi_{\mathrm{ML}}$ violates $\gamma$-PBR using the profile described earlier with $M=3$. The aggregated preference function ${\rm{P}}^{\sigma}$ can be represented by the following matrix:
${\rm{P}}^{\sigma}=\begin{bmatrix}\frac{1}{2}&\frac{1}{2}+\epsilon&\frac{1+\epsilon}{2}\\ \frac{1}{2}-\epsilon&\frac{1}{2}&\frac{1-\epsilon}{2}\\ \frac{1-\epsilon}{2}&\frac{1+\epsilon}{2}&\frac{1}{2}\end{bmatrix}.$ (27)
Now, suppose that group $G_{3}$ manipulates their sub-profile from ${\rm{P}}^{\sigma}_{3}(y_{1}\succ y_{2})=\frac{1}{2}$ to ${\rm{P}}^{\sigma^{\prime}}_{3}(y_{1}\succ y_{2})=0$. Then, the resulting manipulated aggregated preference function ${\rm{P}}^{\sigma^{\prime}}$ is calculated as:
${\rm{P}}^{\sigma^{\prime}}=\begin{bmatrix}\frac{1}{2}&\frac{1}{3}+\epsilon&\frac{1+\epsilon}{2}\\ \frac{2}{3}-\epsilon&\frac{1}{2}&\frac{1-\epsilon}{2}\\ \frac{1-\epsilon}{2}&\frac{1+\epsilon}{2}&\frac{1}{2}\end{bmatrix}.$ (28)
$\Phi_{\mathrm{ML}}$ yields a stochastic policy that depends on the value of $\epsilon$. For example, if $\epsilon=1/12$, the resulting policy is $\pi=\left[\frac{1}{4},\frac{1}{4},\frac{1}{2}\right]$. However, as $\epsilon$ approaches 00, the resulting policy converges to $[0,0,1]$. This shows that $\pi^{\prime}(y_{3})\to 1$ while $w^{\sigma}_{3}=1/3$, thus demonstrating that there exists no $\gamma_{1}>0$ for which $\Phi_{\mathrm{ML}}$ satisfies $\gamma$-PBR.
To show that $\Phi_{\mathrm{MB}}$ violates $\gamma$-PBR, consider the case with $M=3$ where the profile $\sigma$ consists of the following three groups of evaluators:
$\sigma=\{(y_{1}\succ y_{2}\succ y_{3})\times 0.30,\;(y_{2}\succ y_{1}\succ y_{3})\times 0.45,\;(y_{3}\succ y_{1}\succ y_{2})\times 0.25\},$ (29)
where $(y_{1}\succ y_{2}\succ y_{3})$ represents a ranking $r$ and “×$\times 0.30$” indicates that $\sigma_{r}=0.30$. Then, the Borda scores are calculated as $B=[1.3,1.20,0.5]$. Thus, $\Phi_{\mathrm{MB}}(\sigma)=\pi$, where $\pi(y_{1})=1$. Next, suppose the second group strategically misreports their preference from $(y_{2}\succ y_{1}\succ y_{3})$ to $(y_{2}\succ y_{3}\succ y_{1})$. Then, the Borda scores are calculated as $B^{\prime}=[0.85,1.2,0.95]$. The resulting policy is then $\pi^{\prime}(y_{2})=1$, with the population share of the second group being $w^{\sigma}_{2}=0.45$. This example demonstrates that there exists no $\gamma_{1}>0$ for which $\Phi_{\mathrm{MB}}$ satisfies $\gamma$-PBR.
Next, we show that $\Phi_{\mathrm{RD}}$ satisfies all four axioms. $\Phi_{\mathrm{RD}}$ satisfies monotonicity because improving ranking of $y$ cannot decrease the number of evaluators whose top choice is $y$. In addition, $\Phi_{\mathrm{RD}}$ satisfies Pareto efficiency because if $r(y_{j})<r(y_{k})$ for all $r\in S$ with $\sigma_{r}>0$, then we have $w^{\sigma}_{k}=0$ and $\Phi_{\mathrm{RD}}(\sigma)(y_{k})=0$. Additionally, $\Phi_{\mathrm{RD}}$ satisfies $\alpha$-PPR with $\alpha(\sigma)=1$ for all $\sigma$ by its definition, and also satisfy $\gamma$-PBR with $(\gamma_{1},\gamma_{2})=(1,0)$ because each group $G_{k}$ cannot increase $w^{\sigma}_{k}$ by manipulation.
Appendix D Proof of the non-implementability of $\Phi_{\mathrm{RD}}$
Suppose that $\Phi_{\mathrm{RD}}$ can be implemented by a preference learning algorithm $\Psi_{\mathrm{RD}}$. Let $M=3$, and consider two preference profiles, $\sigma_{1}$ and $\sigma_{2}$, defined as follows:
$\begin{split}\sigma_{1}&=\{(y_{1}\succ y_{2}\succ y_{3})\times 1/3,\;(y_{2}\succ y_{1}\succ y_{3})\times 1/3,\;(y_{3}\succ y_{1}\succ y_{2})\times 1/3\},\\ \sigma_{2}&=\{(y_{1}\succ y_{2}\succ y_{3})\times 2/3,\;(y_{3}\succ y_{2}\succ y_{1})\times 1/3\}.\end{split}$ (30)
Both of these profiles induce the same aggregated preference function ${\rm{P}}^{\sigma}={\rm{P}}^{\sigma_{1}}={\rm{P}}^{\sigma_{2}}$, where
${\rm{P}}^{\sigma}=\begin{bmatrix}\frac{1}{2}&\frac{2}{3}&\frac{2}{3}\\ \frac{1}{3}&\frac{1}{2}&\frac{2}{3}\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{2}\end{bmatrix}.$ (31)
Therefore, the preference learning algorithm $\Psi_{\mathrm{RD}}$ would produce the same policy for both $\sigma_{1}$ and $\sigma_{2}$. However, according to the definition of $\Phi_{\mathrm{RD}}$, we have $\Phi_{\mathrm{RD}}(\sigma_{1})=[1/3,1/3,1/3]$ and $\Phi_{\mathrm{RD}}(\sigma_{2})=[2/3,0,1/3]$, which are different policies. This implies that $\Psi_{\mathrm{RD}}$ does not implement $\Phi_{\mathrm{RD}}$, which contradicts our initial assumption. Therefore, $\Phi_{\mathrm{RD}}$ is not implementable by a preference learning algorithm.
Appendix E Proof of Proposition 4.2
First, consider any feasible population share $w$ given a preference function ${\rm{P}}$. By Definition 4.1, there exists a profile $\sigma$ such that $w=w^{\sigma}$ and ${\rm{P}}={\rm{P}}^{\sigma}$. Then, the group-specific preference functions $({\rm{P}}^{\sigma}_{1},\ldots,{\rm{P}}^{\sigma}_{M})$ that constitute ${\rm{P}}^{\sigma}$, satisfy the condition in Equation (3), which implies that $w\in\mathcal{W}({\rm{P}})$.
Next, consider any $w\in\mathcal{W}({\rm{P}})$. By the definition of $\mathcal{W}({\rm{P}})$, there exist $({\rm{P}}_{1},\ldots,{\rm{P}}_{M})\in\mathcal{P}^{M}$ such that ${\rm{P}}=\sum_{k=1}^{M}w_{k}{\rm{P}}_{k}$, where each ${\rm{P}}_{k}$ satisfies ${\rm{P}}_{k}(y_{k}\succ y)=1$ for all $y\neq y_{k}$. Since ${\rm{P}}_{k}\in\mathcal{P}$, there exists a profile $\sigma_{k}$ that induces ${\rm{P}}_{k}$, such that ${\rm{P}}_{k}={\rm{P}}^{\sigma_{k}}$. Now, if we consider an aggregated profile $\sigma:=\sum_{k=1}^{M}w_{k}\sigma_{k}$ by combining these group profiles with the corresponding weights, then the preference function of $\sigma$ will be ${\rm{P}}^{\sigma}=\sum_{k=1}^{M}w_{k}{\rm{P}}^{\sigma_{k}}=\sum_{k=1}^{M}w_{k}{\rm{P}}_{k}={\rm{P}}$ and also $w^{\sigma}=w$. Therefore, $w$ is a feasible population distribution given ${\rm{P}}$.
Appendix F Proof of Theorem 4.4
Consider any $w\in\mathcal{W}({\rm{P}})$. Then, there exists $({\rm{P}}_{1},\ldots,{\rm{P}}_{M})\in\mathcal{P}^{M}$ such that ${\rm{P}}=\sum_{k=1}^{M}w_{k}\,{\rm{P}}_{k}$. Fix an index $i\in[M]$. For any $y\in\mathcal{Y}\setminus\{y_{i}\}$, we have
${\rm{P}}(y_{i}\succ y)=w_{i}\,{\rm{P}}_{i}(y_{i}\succ y)+\sum_{k\neq i}w_{k}\,{\rm{P}}_{k}(y_{i}\succ y)\geq w_{i}$ (32)
since ${\rm{P}}_{i}(y_{i}\succ y)=1$. Taking the minimum over $y\in\mathcal{Y}\setminus\{y_{i}\}$ yields
$\min_{y\in\mathcal{Y}\setminus\{y_{i}\}}{\rm{P}}(y_{i}\succ y)\geq w_{i},$ (33)
which implies $w_{i}\leq u_{i}\;\forall i\in[M]$ and $w\in\overline{\mathcal{W}}({\rm{P}})$. Therefore, $\mathcal{W}({\rm{P}})\subseteq\overline{\mathcal{W}}({\rm{P}})$.
Appendix G Proof of $\mathcal{W}({\rm{P}})=\overline{\mathcal{W}}({\rm{P}})$ under extended preference profiles
Consider an extended profile setting in which each group $G_{k}$ is allowed to provide pairwise comparison data according to its own preference function ${\rm{P}}_{k}$, subject only to the skew-symmetry constraint ${\rm{P}}_{k}(y_{i}\succ y_{j})+{\rm{P}}_{k}(y_{j}\succ y_{i})=1$ for all $y_{i},y_{j}\in\mathcal{Y}$, and the unanimity constraint ${\rm{P}}_{k}(y_{k}\succ y)=1$ for all $y\in\mathcal{Y}$. In this case, the set $\mathcal{P}$ is defined as
$\mathcal{P}:=\left\{{\rm{P}}\mid{\rm{P}}(y\succ y^{\prime})+{\rm{P}}(y^{\prime}\succ y)=1\;\forall y,y^{\prime}\in\mathcal{Y}\right\}.$ (34)
We show that $\overline{\mathcal{W}}({\rm{P}})\subseteq\mathcal{W}({\rm{P}})$ also holds under this setting. Consider any $w\in\overline{\mathcal{W}}({\rm{P}})$. By assumption, $w$ satisfies $w_{i}\leq u_{i}=\min_{y\in\mathcal{Y}\setminus\{y_{i}\}}{\rm{P}}(y_{i}\succ y)$ for all $i\in[M]$. Define each element of $({\rm{P}}_{1},\ldots,{\rm{P}}_{M})$ as
${\rm{P}}_{k}(y_{i}\succ y_{j})=\frac{{\rm{P}}(y_{i}\succ y_{j})-w_{i}}{1-w_{i}-w_{j}}$ (35)
for any $i,j\neq k$, and let ${\rm{P}}_{k}(y_{k}\succ y)=1$, ${\rm{P}}_{k}(y\succ y_{k})=0$ for all $y\in\mathcal{Y}$. Then ${\rm{P}}_{k}(y_{i}\succ y_{j})\in[0,1]$ holds because ${\rm{P}}(y_{i}\succ y_{j})\in[w_{i},1-w_{j}]$ by assumption. The skew-symmetry condition ${\rm{P}}_{k}(y_{i}\succ y_{j})+{\rm{P}}_{k}(y_{j}\succ y_{i})=1$ is also satisfied. Thus, ${\rm{P}}_{k}\in\mathcal{P}$ and ${\rm{P}}_{k}$ can be induced by some profile. Finally, the constraint ${\rm{P}}=\sum_{k=1}^{M}w_{k}{\rm{P}}_{k}$ also holds. Therefore, $w\in\mathcal{W}({\rm{P}})$, implying $\overline{\mathcal{W}}({\rm{P}})\subseteq\mathcal{W}({\rm{P}})$, and hence $\mathcal{W}({\rm{P}})=\overline{\mathcal{W}}({\rm{P}})$.
Appendix H Proof of Theorem 4.6
We first prove monotonicity. Improving the ranking of $y_{i}$ for some evaluator can only increase ${\rm{P}}^{\sigma}(y_{i}\succ y)$ for any $y\neq y_{i}$, and decrease ${\rm{P}}^{\sigma}(y\succ y_{i})$. This implies that $u_{i}$ cannot decrease, while $u_{j}$ for $j\neq i$ cannot increase. Therefore, $\pi(y_{i})=u_{i}/(\sum_{j=1}^{M}u_{j})$ cannot decrease, establishing monotonicity.
Next, we prove Pareto efficiency. Suppose $y_{i}$ is ranked above $y_{j}$ in every input ranking, i.e., $r(y_{i})<r(y_{j})$ for all $r\in S$ with $\sigma_{r}>0$. Then, we have ${\rm{P}}^{\sigma}(y_{i}\succ y_{j})=\sum_{r\in S}\sigma_{r}\cdot\mathbf{1}_{\{r(y_{i})<r(y_{j})\}}=1$ and also ${\rm{P}}^{\sigma}(y_{j}\succ y_{i})=0$. Thus, we get $u_{j}=\min_{y\in Y\setminus\{y_{j}\}}{\rm{P}}^{\sigma}(y_{j}\succ y)=0$. Thus, the resulting policy satisfies $\Phi^{*}(\sigma)(y_{j})=u_{j}/(\sum_{k=1}^{M}u_{k})=0$. Therefore, $\Phi^{*}(\sigma)(y_{i})\geq\Phi^{*}(\sigma)(y_{j})$, establishing that $\Phi^{*}$ satisfies Pareto efficiency.
Appendix I Proof of Lemma 4.7 and Lemma 4.9
Lemma 4.7 follows directly from the fact that $w^{\sigma}_{i}\leq u_{i}$ for all $i\in[M]$, which gives
$\frac{\pi(y_{i})}{w_{i}^{\sigma}}=\frac{u_{i}}{w_{i}^{\sigma}\sum_{j=1}^{M}u_{j}}\geq\frac{1}{\sum_{j=1}^{M}u_{j}}.$ (36)
Next, we show Lemma 4.9. Let $I\subseteq[M]$ be the set of indexes for $\delta$-dominated alternatives, where $|I|=N_{\delta}^{\sigma}$. Then, for any $i\in I$, we have
$u_{i}=\min_{j\in[M]\setminus\{i\}}{\rm{P}}^{\sigma}(y_{i}\succ y_{j})\leq{\rm{P}}^{\sigma}(y_{i}\succ y^{\prime})\leq 1-\delta,$ (37)
where $y^{\prime}_{i}$ denotes an alternative that $\delta$-dominates $y_{i}$. Additionally, let $k\in\operatorname*{arg\,max}_{i\in[M]}w_{i}^{\sigma}$. Then, for any $i\neq k$, we have
$u_{i}=\min_{j\in[M]\setminus\{i\}}{\rm{P}}^{\sigma}(y_{i}\succ y_{j})\leq{\rm{P}}^{\sigma}(y_{i}\succ y_{k})\leq 1-w_{k}^{\sigma}$ (38)
Similarly, let $l\in\operatorname*{arg\,max}_{i\in[M],i\neq k}w_{i}^{\sigma}$, then we have $u_{k}\leq 1-w_{l}^{\sigma}$. Combining these reesults,
$\sum_{i=1}^{M}u_{i}=u_{k}+\sum_{i\neq k,i\notin I}u_{i}+\sum_{i\in I}u_{i}\leq(1-w_{l}^{\sigma})+(N^{\sigma}_{\delta}-1)(1-w_{k}^{\sigma})+(M-N^{\sigma}_{\delta})(1-\delta).$ (39)
In addition, since $u_{i}\geq w_{i}^{\sigma}$, we have $\sum_{i=1}^{M}u_{i}\geq\sum_{i=1}^{M}w_{i}^{\sigma}=1$. Combining both inequalities and plugging $(w_{k}^{\sigma},w_{l}^{\sigma})=(w^{\sigma,1},w^{\sigma,2})$ in, we get the result of Lemmma 4.9 as follows:
[…]
Appendix J Proof of Theorem 4.11
Let $\sigma^{\prime}$ be the profile manipulated by group $G_{k}$, and let $\pi^{\prime}=\Phi^{*}(\sigma^{\prime})$ be the resulting policy. $G_{k}$ aims to maximize
$\pi^{\prime}(y_{k})=\frac{u^{\prime}_{k}}{u^{\prime}_{k}+\sum_{i\neq k}u^{\prime}_{i}},$ (41)
where $u^{\prime}$ represents the value of $u$ after the manipulation. To maximize $\pi^{\prime}(y_{k})$, $G_{k}$ will attempt to maximize $u^{\prime}_{k}$ and minimize $\sum_{i\neq k}u^{\prime}_{i}$. Since increasing the ranking of $y_{k}$ in their profile increases (or at least does not decrease) the value of $u^{\prime}_{k}$ without increasing the value of $\sum_{i\neq k}u^{\prime}_{i}$, the optimal strategy for $G_{k}$ is to truthfully report $y_{k}$ as its top choice. In this strategy, we have $u^{\prime}_{k}=u_{k}$ and the sum $\sum_{i\neq k}u^{\prime}_{i}$ has the following lower bound:
$\sum_{i\neq k}u^{\prime}_{i}\geq\sum_{i\neq k}w^{\sigma}_{i}=1-w^{\sigma}_{k}.$ (42)
Substituting this lower bound into equation (41), we obtain
$\pi^{\prime}(y_{k})\leq\frac{u_{k}}{u_{k}+1-w^{\sigma}_{k}}\leq\frac{1}{2}(w_{k}^{\sigma}+1),$ (43)
where the final inequality holds if $(w_{k}^{\sigma}-u_{k}+1)(w_{k}^{\sigma}-1)\leq 0$, which follows from the fact that $w_{k}^{\sigma},u_{k}\in[0,1]$.
Appendix K Population-bounded strategyproofness guarantee
In social choice theory, a mechanism is considered strategyproof if participants cannot benefit (i.e., increase their utility) by misreporting their true preferences [36], regardless of what other participants report. In our preference learning framework, we assume each group $G_{k}$’s utility is the probability assigned to its top choice, represented by $\pi(y_{k})$. A preference learning algorithm is strategyproof if no group can improve its outcome by misreporting preferences. However, as noted by Buening et al. [21], strict strategyproofness is typically too restrictive and is not satisfied by the conventional preference learning algorithms. Our method does not satisfy strict strategyproofness like other methods, but satisfies a weaker form that provides a bounded guarantee on the maximum potential gain from strategic misreporting in equilibrium.
Let $\sigma^{\prime}$ denote the profile resulting from strategical misreporting by all groups, and let $\pi^{\prime}=\Phi^{*}(\sigma^{\prime})$ be the resulting policy. Each group $G_{k}$ aims to maximize $\pi^{\prime}(y_{k})$, which involves maximizing $u^{\prime}_{k}$ and minimizing $\sum_{i\neq k}u^{\prime}_{i}$.
Since improving the ranking of $y_{k}$ in their reported preferences increases (or at worst, does not decrease) the value of $u^{\prime}_{k}$ without increasing $\sum_{i\neq k}u^{\prime}_{i}$, the optimal strategy for $G_{k}$ is to truthfully report $y_{k}$ as their top choice. Hence, all groups truthfully report their top choice regardless of other groups’ strategies, meaning ${\rm{P}}^{\prime}_{k}(y_{k}\succ y)=1$ for all $y\neq y_{k}$, where ${\rm{P}}^{\prime}_{k}$ denotes the reported preference function of $G_{k}$.
In this equilibrium, following steps analogous to the proof of Theorem 4.11, we have:
$\pi^{\prime}(y_{k})\leq\frac{u^{\prime}_{k}}{u^{\prime}_{k}+1-w^{\sigma}_{k}}\leq\frac{1}{2}(w_{k}^{\sigma}+1)\quad\forall k\in[M].$ (44)
Note that $\gamma(w_{k}^{\sigma})$ is not a tight bound. Further exploration into tighter bounds and detailed analysis of each group’s strategic behavior is left for future research.
Appendix L Proof of Proposition 4.13
Suppose $M=2$ and ${\rm{P}}^{\sigma}(y_{1}\succ y_{2})\in(0.5,1)$, so $y_{1}$ is the Condorcet winner. If a PSCF $\Phi$ satisfies Condorcet consistency, it must return the deterministic policy $\pi(y_{1})=1$. However, this violates the $\alpha$-PPR axiom because $\pi(y_{2})=0$ while $w^{\sigma}_{2}>0$, which implies that $\pi(y_{2})/w^{\sigma}_{2}=0$ cannot be lower bounded by any $\alpha(\sigma)>0$.
Appendix M Proof of Proposition 4.14
Suppose $y_{i}$ is a Condorcet winner. Then ${\rm{P}}^{\sigma}(y_{i}\succ y_{j})>0.5$ for all $j\neq i$, which implies that $u_{i}>0.5$. For any other $j\neq i$, we have $u_{j}\leq{\rm{P}}^{\sigma}(y_{j}\succ y_{i})<0.5$. Therefore, $y_{i}$ has the highest $u_{i}$, i.e., $i\in\operatorname*{arg\,max}_{j\in[M]}u_{j}$, and $\Phi^{\infty}$ returns $\pi(y_{i})=1$, satisfying Condorcet consistency.
Appendix N Finite Behavior of $\Phi^{\beta}$
The following proposition quantifies how large the parameter $\beta$ needs to be to ensure that a Condorcet winner receives a sufficiently high probability under the softmax policy.
Proposition N.1 (Condorcet consistency at finite $\beta$).
Let $y_{i}$ be a Condorcet winner with $u_{i}>0.5$. Then, the softmax policy satisfies $\pi(y_{i})\geq\alpha_{c}$ if
$\beta\geq\frac{1}{u_{i}-0.5}\log\left(\frac{(M-1)\alpha_{c}}{2(1-\alpha_{c})}\right).$ (45)
Proof.
Since $y_{i}$ is a Condorcet winner, we have $u_{j}\leq{\rm{P}}^{\sigma}(y_{j}\succ y_{i})=1-{\rm{P}}^{\sigma}(y_{i}\succ y_{j})<0.5$ for any $j\neq i$. From the given condition
$\beta\geq\frac{1}{u_{i}-0.5}\log\left(\frac{(M-1)\alpha_{c}}{2(1-\alpha_{c})}\right),$ (46)
we can establish the following lower bound:
$u_{i}\exp{(\beta u_{i})}\geq\frac{\alpha_{c}}{1-\alpha_{c}}(M-1)(0.5\exp{(0.5\beta))}\geq\frac{\alpha_{c}}{1-\alpha_{c}}\sum_{j\neq i}u_{j}\exp{(\beta u_{j})}.$ (47)
Thus, the softamx policy satisfies
$\pi(y_{i})=\frac{u_{i}\exp{(\beta u_{i})}}{u_{i}\exp{(\beta u_{i})}+\sum_{j\neq i}u_{j}\exp{(\beta u_{j})}}\geq\alpha_{c}.$ (48)
In addition, it can be shown that the $\alpha$-PPR guarantee deteriorates as $\beta\to\infty$, since the lower bound in Lemma 4.7 becomes
$\frac{\pi(y_{i})}{w_{i}^{\sigma}}\geq\left(\sum_{j=1}^{M}u_{j}\exp\left(\beta(u_{j}-u_{i})\right)\right)^{-1},$ (49)
which converges to zero as $\beta\to\infty$, unless $u_{i}=\max_{j\in[M]}u_{j}$.
Appendix O Proof of Proposition 4.17
First, we show that $\Phi^{\beta}$ violates PMC for any $\beta\geq 0$. It suffices to demonstrate that there exists a profile $\sigma$ with a PMC ranking $r^{\sigma}$ for which $\Phi^{\beta}(\sigma)(y_{i})<\Phi^{\beta}(\sigma)(y_{j})$ while $r^{\sigma}(y_{i})<r^{\sigma}(y_{j})$.
Consider the following profile $\sigma$ with $M=3$:
$\sigma=\{(y_{1}\succ y_{2}\succ y_{3})\times 0.3,\;(y_{2}\succ y_{3}\succ y_{1})\times 0.1,\;(y_{3}\succ y_{1}\succ y_{2})\times 0.3,\;(y_{3}\succ y_{2}\succ y_{1})\times 0.3\},$ (50)
which yields the following preference function:
${\rm{P}}^{\sigma}=\begin{bmatrix}0.5&0.6&0.3\\ 0.4&0.5&0.4\\ 0.7&0.6&0.5\end{bmatrix}.$ (51)
Then, the PMC ranking satisfies $r^{\sigma}(y_{3})<r^{\sigma}(y_{1})<r^{\sigma}(y_{2})$, as ${\rm{P}}^{\sigma}(y_{3}\succ y_{1}),{\rm{P}}^{\sigma}(y_{3}\succ y_{2}),{\rm{P}}^{\sigma}(y_{1}\succ y_{2})>0.5$. However, we have $u_{1}=0.3$ and $u_{2}=0.4$. Since $u_{1}<u_{2}$, it follows that $\Phi^{\beta}(\sigma)(y_{1})<\Phi^{\beta}(\sigma)(y_{2})$ regardless of $\beta$, contradicting $r^{\sigma}(y_{1})<r^{\sigma}(y_{2})$. Therefore, $\Phi^{\beta}$ violates PMC regardless of $\beta$.
Next, we show that $\Phi^{\infty}$ satisfies PMC. Consider a profile $\sigma$ with its PMC ranking $r^{\sigma}$, and let $y^{*}\in\mathcal{Y}$ be the alternative ranked first in $r^{\sigma}$ (i.e., $r^{\sigma}(y^{*})=1$). Then, $y^{*}$ must be a Condorcet winner, as ${\rm{P}}^{\sigma}(y^{*}\succ y)>1/2$ for all $y\neq y^{*}$. Thus, $\pi:=\Phi^{\infty}(\sigma)$ is a deterministic policy with $\pi(y^{*})=1$. Consequently, we have $\pi(y^{*})>\pi(y)$ for all $y\neq y^{*}$ and trivially $\pi(y_{i})=\pi(y_{j})=0$ for any $y_{i},y_{j}\neq y^{*}$, satisfying the condition for PMC.
Appendix P Derivation of the loss function in the initial phase
Step 1: LP reformulation
Recall the definition $u(y\mid x):=\min_{z\in\mathcal{Y}\setminus\{y\}}{\rm{P}}(y\succ z\mid x)$. Each $u(y\mid x)$ can be rewritten as
$u(y\mid x)=\sum_{z\in\mathcal{Y}}{\rm{P}}(y\succ z\mid x)\mu^{*}(z\mid x,y),$ (52)
where a selector distribution $\mu^{*}(\,\cdot\mid x,y)\in\Delta(\mathcal{Y})$ places all its mass on the minimizer of ${\rm{P}}(y\succ\cdot\mid x)$. Such $\mu^{*}$ and the corresponding pointwise minimum can be obtained via the following linear programming (LP):
$u(y\mid x)=\min_{\mu(\cdot\mid x,y)\in\Delta(\mathcal{Y})}\sum_{z\in\mathcal{Y}}{\rm{P}}(y\succ z\mid x)\mu(z\mid x,y).$ (53)
Step 2: Aaggregation of pointwise LPs
Assume the data-generating distribution $\rho(\cdot)$ and $\pi_{d}(\;\cdot\mid x)$ have full support. Multiplying and dividing (53) by $\pi_{d}(z\mid x)$ and then taking the expectation over all $z\in\mathcal{Y}$ gives
$u(y\mid x)=\min_{\mu(\cdot\mid x,y)\in\Delta(\mathcal{Y})}{\mathbb{E}}_{z\sim\pi_{d}(\cdot\mid x)}\left[\frac{{\rm{P}}(y\succ z\mid x)\mu(z\mid x,y)}{\pi_{d}(z\mid x)}\right].$ (54)
Next, we aggregate these pointwise LPs by multiplying each pointwise objective by $\rho(x)\pi_{d}(y\mid x)$ and summing over all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. We also add the symmetrical term with swapped $y$ and $z$, which does not change the optimal solution:
$\mu^{*}\in\operatorname*{arg\,min}_{\mu:\mathcal{X}\times\mathcal{Y}\mapsto\Delta(\mathcal{Y})}\mathbb{E}_{(x,y,z)\sim(\rho,\pi_{d},\pi_{d})}\left[\frac{{\rm{P}}(y\succ z\mid x)\mu(z\mid x,y)}{\pi_{d}(z\mid x)}+\frac{{\rm{P}}(z\succ y\mid x)\mu(y\mid x,z)}{\pi_{d}(y\mid x)}\right].$ (55)
Step 3: Empirical counterpart
Given an offline preference dataset $\mathcal{D}$, we approximate the expectation in (55) using its empirical counterpart and restrict the function class to a parameterized family ${\mathcal{F}}_{\mu}$:
$\hat{\mu}\in\operatorname*{arg\,min}_{\mu\in{\mathcal{F}}_{\mu}}\frac{1}{N}\sum_{i=1}^{N}\frac{\mu(y^{l}_{i}\mid x_{i},y^{w}_{i})}{\pi_{d}(y^{l}_{i}\mid x_{i})}.$ (56)
Given the estimate $\hat{\mu}$, we can estimate $\hat{u}$ using the definition (53) with the estimated preference function $\hat{{\rm{P}}}$:
$\hat{{\rm{P}}}(y\succ z\mid x):=\begin{cases}\frac{N(x,y,z)}{N(x,y,z)+N(x,z,y)}&\text{if }N(x,y,z)+N(x,z,y)>0,\\ 1/2,&\text{otherwise},\end{cases}$ (57)
where $N(x,y,z):=|\{i\in[N]\mid(x_{i},y_{i}^{w},y_{i}^{l})=(x,y,z)\}|$.
Appendix Q Practical techniques for LLM fine-tuning
Loss function with KL divergence
Substituting the KL divergence into $L^{\pi}$ from equation (12), the loss function becomes
$\begin{split}&{\mathbb{E}}_{x\sim\rho}\left[D_{\mathrm{KL}}\left(\pi(\cdot\mid x)\bigg{\|}\sum_{z\in\mathcal{Y}}\hat{P}(\cdot\succ z\mid x)\hat{\mu}(z\mid x,\cdot)\right)\right]\\ &={\mathbb{E}}_{(x,y)\sim(\rho,\pi_{d})}\left[\frac{\pi(y\mid x)}{\pi_{d}(y\mid x)}\log\frac{\pi(y\mid x)}{\sum_{z\in\mathcal{Y}}\hat{P}(y\succ z\mid x)\hat{\mu}(z\mid x,y)}\right].\end{split}$ (58)
Using the offline dataset and function approximation, we obtain
$\begin{split}\hat{\pi}\in\operatorname*{arg\,min}_{\pi\in{\mathcal{F}}_{\pi}}\;&{\mathbb{E}}_{(x,y^{w},y^{l})\sim D}\Bigg{[}\frac{\pi(y^{w}\mid x)}{\pi_{d}(y^{w}\mid x)}\log\frac{\pi(y^{w}\mid x)}{(\frac{1}{2}+\frac{1}{2}\hat{\mu}(y^{w}\mid x,y^{l}))\exp{(\frac{\beta}{2}+\frac{\beta}{2}\hat{\mu}(y^{w}\mid x,y^{l}))}}\\ &\quad\quad\quad\quad\quad\quad\quad+\frac{\pi(y^{l}\mid x)}{\pi_{d}(y^{l}\mid x)}\log\frac{\pi(y^{l}\mid x)}{(\frac{1}{2}-\frac{1}{2}\hat{\mu}(y^{l}\mid x,y^{w}))\exp{(\frac{\beta}{2}-\frac{\beta}{2}\hat{\mu}(y^{w}\mid x,y^{l}))}}\Bigg{]}.\\ \end{split}$ (59)
Regularization via reference policy
Fine-tuning large language models (LLMs) requires maintaining alignment with a reference policy, typically the pretrained model. To prevent excessive drift, we incorporate KL-divergence regularization terms into the training objectives for both the selector model $\mu$ and the policy model $\pi$. Specifically, we add the following regularization terms to the loss functions in Phase 1 and Phase 2:
$\beta_{\mu}{\mathbb{E}}_{(x,y)\sim(\rho,\pi_{d})}\left[D_{\mathrm{KL}}\left(\mu(\cdot\mid x,y)\bigg{\|}\pi_{\mathrm{ref}}(\cdot\mid x,y)\right)\right],\quad\beta_{\pi}{\mathbb{E}}_{x\sim\rho}\left[D_{\mathrm{KL}}\left(\pi(\cdot\mid x)\bigg{\|}\pi_{\mathrm{ref}}(\cdot\mid x)\right)\right]$ (60)
Training with single model
To reduce computational cost, we propose to train both $\mu$ and $\pi$ using a single model. This is enabled by encoding structural differences through specialized input formats. Specifically, $\pi(\;\cdot\mid x)$ selects preferred responses given a prompt, while $\mu(\;\cdot\mid x,y)$ selects responses given a prompt and a candidate response. By distinguishing these cases with separator tokens, we achieve performance comparable to training separate models, while improving memory usage and training efficiency.
Appendix R Additional details of experiments
Datasets and models
The following datasets and models are used in the experiments:
Dataset generation
We considered two groups of evaluators, $G_{1}$ and $G_{2}$, defined across three categories: expertise, informativeness, and style. For each of the 10,000 instruction samples from the Alpaca-GPT4 dataset [32], group-specific responses were generated using the Llama-3.1-8B-Instruct model [34], following the prompts listed in Table 2.
Table 2: Prompts used for generating responses from each group
Category Prompt
Expertise (1) Generate a response that can be easily understood by an elementary school student.
(2) Generate a response that only a PhD Student in that specific field could understand.
Informativeness (1) Generate a response that is concise and to the point, without being verbose.
(2) Generate a response that is very informative, without missing any background information.
Style (1) Generate a response that is friendly, witty, funny, and humorous, like a close friend.
(2) Generate a response that answers in an unfriendly manner.
Training hyperparameters
We fine-tuned the Qwen2-1.5B-Instruct model [31] using each algorithm, where both the reference policy $\pi_{\mathrm{ref}}$ and the data sampling policy $\pi_{d}$ are set to the same base model. All algorithms are trained on the same offline dataset for the same number of iterations. NLHF is not included in the comparison, as their algorithm does not support offline learning. Specific training hyperparameters are provided in Table 3. Each training run utilizes one H100 GPU, requiring approximately one hour per epoch with about 40GB of memory usage (using LoRA and mixed precision training (FP16)).
Table 3: Training hyperparameters
Hyperparameter Value
Training & Reference Model Qwen2-1.5B-Instruct
Learning Rate 1e-4
Batch Size 8
Epochs 3
Optimizer AdamW
Gradient Clipping 1.0
Learning Rate Scheduler Linear
Warmup Steps 100
KL Coefficient 0.02
LoRA Rank 32
LoRA $\alpha$ 16
Evaluation metrics
We evaluate the fine-tuned policy using two key metrics. First, the win rate against the reference policy is evaluated. For each data point, one response from the trained policy and another from the reference policy is sampled. Then, the annotation model (Llama-3.1-8B-Instruct) is used to label which response is preferred by a evaluator from a group ($G_{1}$ or $G_{2}$) randomly selected according to the population shares $(w_{1}^{\sigma},w_{2}^{\sigma})$. The win rate is the percentage of response from the trained policy is preferred over the reference policy’s response, computed over all $10,000$ samples.
Next, the PPR level is evaluated. We sample a response from the trained policy and use the annotation model to classify its group. Table 4 shows the prompts used to classify the group of generated responses. Based on these classifications, we estimate the policy’s PPR level, computed as $\alpha(\sigma)=\min_{i\in\{1,2\}}\pi(y_{i})/w^{\sigma}_{i}$. This metric quantifies how well the policy’s output distribution aligns with the true population distribution, where a higher $\alpha(\sigma)$ indicates better population-proportional representation.
Table 4: Prompts used for classification
Category Prompt
Expertise Does the expertise level of this response align more closely with the elementary level or the PhD student level? Please answer with only one of these exact options: ‘elementary’ or ‘PhD’.
Informativeness Is this response concise (but less informative) or informative (but less concise)? Please answer with only one of these exact options: ‘concise’ or ‘informative’.
Style Is this response friendly or unfriendly? Please answer with only one of these exact options: ‘friendly’ or ‘unfriendly’.
Appendix S Broader impacts
This paper introduces a novel preference learning framework that aims to enhance population-proportional representation across diverse preferences, offering the potential for positive societal impact by mitigating biases within AI systems. Nevertheless, similar to any preference learning technique, it carries the risk of being misused to perpetuate existing biases, whether through the utilization of non-representative datasets or through design choices that unintentionally favor particular viewpoints. We recognize these potential concerns and emphasize the importance of thoughtful attention to data collection and algorithm design to promote positive impact and minimize harm.